OpenStudy (anonymous):

I NEED TO GRAPH THIS AND ALSO SHOW THE SOLUTIONS PLEASE HELP! 2x+ y=-3 2x + 4y = 6

OpenStudy (anonymous):

First solve your two equations for y: 2x+y=-3 ; add 3 to both sides and subtract y from both sides 2x+3=-y ;multiply both sides by -1 y=-2x-3 ; equation (1) 2x+4y=6 2x-6=-4y y=(2x-6)/-4 OR y=-2/4x - 6/4 y=-1/2x-3/2 ;equation (2) You know your slopes for each (-2 for (1) and -1/2 for (2)) and the y-intercept. You know its a straight line so just plot the intercept and graph. Hope that helps

OpenStudy (anonymous):

thank you - what do i plug in to graph it?

OpenStudy (anonymous):

No problem, again, see what i wrote at the end of the answer, use the y-intercepts and the slope to draw the line. Plot the intercept first then using the slope (rise/run) draw the line.

OpenStudy (anonymous):

oooh okay, i'm sorry - it took me a minute to figure it out - AWFUL at matjh. i see now and it's plugged in my calc!